# A priori estimates for parabolic Monge-Ampère type equations

### v1

@@ 2.3. Step 3: global $\tilde{C}^{2,\alpha}$ estimates
Recall that $v=T_{k}(u-\phi)$ in $B_{\rho_{0}}(Y_{0})$ with $T_{k}=\partial_{k}+\rho_{k}\partial_{n}$, $1\leq k<n$, where we take $\rho_{0}$ to be small. If we define $\tilde{u}(x,t)=u(x^{\prime},x_{n}+\rho(x^{\prime}),t)$, then $\tilde{u}_{ij}=T_{i}T_{j}u$, for any $1\leq i,j\leq n$. What we have proved above is equivalent to
$||\tilde{u}_{kl}||_{\tilde{C}^{\alpha}(\overline{G})}\leq C\text{ for any }1\leq k<n\text{ and }1\leq l\leq n.$ (2.27)
To prove the $\tilde{C}^{\alpha}$ estimate for $\tilde{u}_{nn}$, note that the equation (2.1) can be written as $\det(P_{i}P_{j}\tilde{u})=\tilde{u}_{t}$, with $P_{k}=\partial_{k}-\rho_{k}\partial_{n}$ for $1\leq k<n$ and $P_{n}=\partial_{n}$, which is equivalently
$Q_{1}\tilde{u}_{nn}+Q_{2}=\tilde{u}_{t},$
where $Q_{1}$ and $Q_{2}$ are polynomials of $\tilde{u}_{kl}$ with $k+l<2n$ and $\rho$ up to the second derivatives. By Theorem 2.4, the eigenvalues of the Hessian $D^{2}u$ have positive lower and upper bounds; therefore $Q_{1}$ has a positive lower bound. This fact together with Theorem 2.6 and (2.27) leads to the $\tilde{C}^{\alpha}$ estimate of $\tilde{u}_{nn}$. From this the $\tilde{C}^{\alpha}$ estimate of $D^{2}u$ in $\Omega\times(\epsilon,T]$ follows by the arbitrariness of $Y_{0}$ and the interior higher regularity of $u$ in Theorem 2.5.
@@ 3. Existence and uniqueness
Equation (2.1) is uniformly parabolic by virtue of Theorem 2.4. Hence with some compatibility conditions on $\phi$, we have the following short-time existence theorem by the inverse function theorem. The proof is a modification of the argument in [12], Theorem 2.5.7.
@@
Suppose that $\Omega$ is a uniformly convex domain with smooth boundary. $\varphi$ is a smooth and uniformly convex function on $\overline{\Omega}$. $h$ is a smooth function on $\partial\Omega\times[0,T]$ with $h_{t}>c_{0}>0$. Suppose also that the equation
@@
$h(0)=\varphi|_{\partial\Omega}\quad\text{ and }\quad h_{t}(0)=\det D^{2}\varphi|_{\partial\Omega}.$
Then for any $0<\alpha\leq 1$, there is a small positive constant $\epsilon$ such that the initial boundary value problem (3.1) has a unique solution $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$, which is uniformly parabolically convex. Here $Q_{2\epsilon}=\Omega\times(0,2\epsilon]$ and $\epsilon<T$ depends on $n,\Omega,\varphi,h$ and $\alpha$.
@@
$g=-w_{t}+\det D^{2}w\in\tilde{C}^{\alpha}(\overline{Q}_{T_{0}}).$
@@
$\Phi(u)=(-u_{t}+\det D^{2}u,~{}~{}u|_{t=0},~{}~{}u|_{\partial\Omega\times[0,T_{0}]}),$ (3.3)
@@
$S=\{(a,b)\in C^{2,\alpha}(\overline{\Omega})\times\tilde{C}^{2,\alpha}(\partial\Omega\times[0,T_{0}]):\;a|_{\partial\Omega}=b(0),\det D^{2}a|_{\partial\Omega}=b_{t}(0)\}.$
@@
$L\eta=(-\eta_{t}+(\det D^{2}w)w^{ij}\eta_{ij},~{}~{}\eta(0),~{}~{}\eta|_{\partial\Omega\times[0,T_{0}]}),\quad\text{for any $\eta\in V$.}$
The construction of $w$ and $S$ ensures the compatibility condition for $L$. Hence by Theorem 5.2 in Chapter 4 of [23], $L$ is an isomorphism from the tangent space of $V$ at $w$ to that of $W$ at $(g,\varphi,h)$. Then by the inverse function theorem, $\Phi$ is an $C^{1}$ diffeomorphism from $B_{\tilde{\rho}}(w)$, a small neighborhood of $w$ in $\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$ to an open neighborhood $U$ of $(g,\varphi,h)$ in $W$. Take $\tilde{\rho}$ small enough, then the functions in $B_{\tilde{\rho}}(w)$ is still uniformly convex in $x$.
@@
Since $g_{\epsilon}=0$ for $0\leq t\leq 2\epsilon$, then $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$ is a uniformly parabolically convex solution to equation (3.1) in $\Omega\times(0,2\epsilon]$. ∎
@@
$u^{(0)}=\varphi,\quad u^{(k+1)}(x)=\frac{\partial^{k}}{\partial t^{k}}\Big{|}_{t=0}\det D^{2}u(t,x).$
Here the right term of the second equality can be recursively expressed by derivatives of $\varphi$.
Lemma 3.3.
Suppose that $u,v\in\tilde{C}^{2}(Q_{T})\cap C(\overline{Q}_{T})$ are uniformly convex in $x$. If
{$-u_{t}+\det D^{2}u&\geq-v_{t}+\det D^{2}v\quad\ \text{ in }Q_{T},\\ u&\leq v\quad\quad\quad\quad\quad\quad\quad\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
Then $u\leq v$ in $Q_{T}$.
Proof.
Note that
00 0 $\geq(u-v)_{t}+\det D^{2}v-\det D^{2}u$
$=(u-v)_{t}+\int_{0}^{1}\frac{d}{ds}\det D^{2}(sv+(1-s)u)\,ds$
$=w_{t}-\bar{a}^{ij}w_{ij},$
where $w=u-v$, and the matrix $(\bar{a}^{ij})$ is
$\int_{0}^{1}\det D^{2}(sv+(1-s)u)(sD^{2}u+(1-s)D^{2}v)^{-1}\,ds.$
Then $(\bar{a}^{ij})>0$. Since $w\leq 0$ on $\partial_{p}Q_{T}$, we have $w\leq 0$ in $Q_{T}$ by comparison theorem of linear parabolic equations. ∎
The uniqueness of solutions to the equation (2.1) then follows immediately from the above lemma.
@@
Let $\Omega$ be a uniformly convex bounded domain with smooth boundary and $\phi$ be a continuous function in $\overline{Q}_{T}$. Then equation (2.1) admits at most one $\tilde{C}^{2}$ smooth and parabolically convex solution.
@@
Additionally, the higher order regularity of $u$ can also be obtained by classic regularity theory of linear parabolic equations iteratively if $\partial\Omega$ and $\phi$ are smooth enough, and equation (2.1) admits such solution uniquely if the equation satisfies higher order compatibility conditions in Remark 3.2. Specifically, let $k\geq 2$, $0<\beta\leq 1$. Assume that $\partial\Omega$ is of class $C^{k+2,\beta}$, $\phi\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$, and equation (2.1) satisfies the compatibility condition of order $[\frac{k+\beta}{2}]+1$. Then the equation admits a unique parabolically convex solution $u\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$. ∎
@@ 4. Gauss curvature flow
{$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{\frac{n+1}{2}}}\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\,\,\text{on }\partial_{p}Q_{T}.\end{aligned}\right.$ (4.1)
@@
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be a parabolically convex solution of equation (4.1). Suppose that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$, $\partial\Omega\in C^{4}$, $\phi\in\tilde{C}^{4}(\overline{Q}_{T})$, $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$ and $\phi(\cdot,0)$ is uniformly convex. Then we have
@@
$\mathcal{L}^{\prime}$ $=-\partial_{t}+\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta}}u^{ij}\partial_{ij}-2\beta\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta+1}}u_{i}\partial_{i}$ (4.2)
@@
$\overline{w}=\tilde{\phi}-\frac{1}{q}\log(1+kd_{\partial\Omega})$ (4.3)
@@
Finally, we take $k$ sufficiently large to satisfy $\log(k\delta)\geq q(\max_{\overline{\Omega}}|\tilde{\phi}|+\max_{\overline{Q}_{T}}|u|)$, then $\overline{w}\leq u$ on $\partial\Omega_{\delta}\times\{t_{0}\}$. Therefore $u\geq\overline{w}$ in $\Omega_{\delta}\times\{t_{0}\}$ by the comparison principle. Then for the exterior normal vector $\nu$ at any point $x_{0}$ on $\partial\Omega$, we have $u_{\nu}\leq w_{\nu}\leq C$. By convexity of $u$ in $x$ and arbitrariness of $t_{0}$, this proves the bound for $|Du|$ on $\overline{Q}_{T}$.
@@
$C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u\leq CI_{n-1}\quad\text{on}\quad\partial_{p}Q_{T}.$ (4.4)
@@
$|\mathcal{L}^{\prime}\mathcal{T}(u-\phi)|\leq\underline{a}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}$ (4.5)
@@
$|\mathcal{T}(u-\phi)|\leq\bar{a}|x^{\prime}|^{2},\quad\text{on $(B_{\varepsilon}\cap\partial\Omega)\times[0,T]$}.$ (4.6)
@@
$A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in $(B_{\varepsilon}\cap\Omega)\times\{0\}$ for all positive constant $A$.}$ (4.7)
@@
$\begin{split}\mathcal{L}^{\prime}\zeta&\geq c_{1}\Big{(}\sum_{i=1}^{n}u^{ii}+Mu^{nn}\Big{)}-c_{2}\\ &\geq\frac{1}{2}c_{1}\sum_{i=1}^{n}u^{ii}-c_{2}+c_{3}M^{1/n}.\end{split}$ (4.8)
@@
$\mathcal{L}^{\prime}(A\zeta\pm\mathcal{T}(u-\phi))\geq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times(0,T].$ (4.9)
@@
$\zeta\leq-\frac{1}{2}\mu\sum_{\beta<n}x_{\beta}^{2}+O(|x^{\prime}|^{3})\leq-\frac{1}{4}\mu|x^{\prime}|^{2},\quad\text{on $(\overline{B_{\varepsilon}}\cap\partial\Omega)\times[0,T]$.}$ (4.10)
@@
$\begin{split}\zeta&\leq\frac{1}{2}(B_{\alpha\beta}-\mu\delta_{\alpha\beta})x_{\alpha}x_{\beta}+\frac{1}{2}Mx_{n}^{2}-x_{n}\\ &\leq-\frac{1}{2}\mu|x^{\prime}|^{2}+C|x^{\prime}|^{3}+Mx_{n}^{2}\\ &\leq-c_{4}\varepsilon^{2},\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$ (4.11)
@@
$\begin{split}\zeta&\leq C|x^{\prime}|^{2}+Mx_{n}^{2}-x_{n}\\ &\leq(M+C)\varepsilon^{2}-c_{5}\varepsilon\\ &\leq-\frac{1}{2}c_{5}\varepsilon,\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$ (4.12)
@@
$A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{on}\quad\partial(B_{\varepsilon}\cap\Omega)\times[0,T],$ (4.13)
@@
$A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times[0,T].$ (4.14)
@@
$1/C\leq\lambda(D^{2}u)\leq C\quad\text{on }\partial_{p}Q_{T}$ (4.16)
@@
Assume $\Omega$ is a smooth and uniformly convex domain in $\mathbb{R}^{n}$, $\phi$ is a smooth function in $\overline{Q}_{T}$ with $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$, and $\phi(\cdot,0)$ is uniformly convex. If equation (4.1) satisfies the compatibility condition of any order, then there exists a unique smooth parabolically solution $u$ to the initial boundary value problem (4.1).

### v2

@@ 2.3. Step 3: global $\tilde{C}^{2,\alpha}$ estimates
Recall that $v(x,t)=T_{k}(u-\phi)(x^{\prime},x_{n}+\rho(x^{\prime}),t)$ in $\Psi(B_{\rho_{0}}(Y_{0}))$ with $T_{k}=\partial_{y_{k}}+\rho_{y_{k}}\partial_{y_{n}}$, $1\leq k<n$, where we take $\rho_{0}>0$ small enough. If we define $\tilde{u}(x,t)=u(x^{\prime},x_{n}+\rho(x^{\prime}),t)$, then $\tilde{u}_{ij}=T_{i}T_{j}u$, for any $1\leq i,j\leq n$. What we have proved above is equivalent to
(2.32) $||\tilde{u}_{kl}||_{\tilde{C}^{\alpha}(\overline{G})}\leq C\text{ for any }1\leq k<n\text{ and }1\leq l\leq n.$
To prove the $\tilde{C}^{\alpha}$ estimate for $\tilde{u}_{nn}$, note that the equation (2.1) can be written as $\det(P_{i}P_{j}\tilde{u})=\tilde{u}_{t}+\psi(\Psi^{-1}(x),t)$, with $P_{k}=\partial_{x_{k}}-\rho_{x_{k}}\partial_{x_{n}}$ for $1\leq k<n$ and $P_{n}=\partial_{x_{n}}$, which is equivalently
$Q_{1}\tilde{u}_{nn}+Q_{2}=\tilde{u}_{t}+\psi(\Psi^{-1}(x),t),$
where $Q_{1}$ and $Q_{2}$ are polynomials of $\tilde{u}_{kl}$ with $k+l<2n$ and $\rho$ up to the second derivatives. By Theorem 2.5, the eigenvalues of the Hessian $D^{2}u$ have positive lower and upper bounds; therefore, $Q_{1}$ has a positive lower bound. This fact together with Theorem 2.7 and (2.32) leads to the $\tilde{C}^{\alpha}$ estimate of $\tilde{u}_{nn}$. From this the $\tilde{C}^{\alpha}$ estimate of $D^{2}u$ in $\Omega\times(\epsilon,T]$ follows by the arbitrariness of $Y_{0}$ and the interior higher regularity of $u$ in Theorem 2.6. ∎
@@ 3. Existence and uniqueness
Equation (2.1) is uniformly parabolic by virtue of Theorem 2.5. Hence with some compatibility conditions on $\phi$, we have the following short-time existence theorem by the inverse function theorem. The proof is a modification of the argument in [9, Theorem 2.5.7].
@@
Let $\Omega$ be a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\varphi$ is a smooth and uniformly convex function on $\overline{\Omega}$, $\psi$ is a smooth function on $\overline{\Omega}\times[0,T]$ and $h$ is a smooth function on $\partial\Omega\times[0,T]$. Suppose also that the equation
@@
$h(0)=\varphi|_{\partial\Omega}\quad\text{ and }\quad-h_{t}(0)+\det D^{2}\varphi|_{\partial\Omega}=\psi(\cdot,0)|_{\partial\Omega}.$
Then for any $0<\alpha\leq 1$, there is a small positive constant $\epsilon$ such that the initial boundary value problem (3.1) has a unique solution $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$, which is uniformly convex in $x$ for all $t\in[0,2\epsilon)$. Here we have $||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})}\leq C$ and $Q_{2\epsilon}=\Omega\times(0,2\epsilon]$. The constants $\epsilon$ and $C$ depend on $n$, $\Omega$, the uniform convexity of $\varphi$, $||\psi||_{\tilde{C}^{1}(\overline{\Omega}\times[0,T])}$, $||\varphi||_{\tilde{C}^{3}(\overline{\Omega})}$, and $||h||_{\tilde{C}^{2}(\partial\Omega\times[0,T])}$.
@@
$g=-w_{t}+\det D^{2}w-\psi\in\tilde{C}^{\alpha}(\overline{Q}_{T_{0}}).$
@@
(3.3) $\Phi(u)=(-u_{t}+\det D^{2}u-\psi,~{}~{}u|_{t=0},~{}~{}u|_{\partial\Omega\times[0,T_{0}]}),$
@@
$S=\{(a,b)\in C^{2,\alpha}(\overline{\Omega})\times\tilde{C}^{2,\alpha}(\partial\Omega\times[0,T_{0}]):\;a|_{\partial\Omega}=b|_{t=0},\det D^{2}a|_{\partial\Omega}=(\psi+b_{t})|_{t=0}\}.$
@@
$L\eta=(-\eta_{t}+(\det D^{2}w)w^{ij}\eta_{ij}-\psi_{t},~{}~{}\eta(0),~{}~{}\eta|_{\partial\Omega\times[0,T_{0}]}),\quad\text{for any $\eta\in V$.}$
The construction of $w$ and $S$ ensures the compatibility condition for $L$. Hence by Theorem 5.2 in Chapter 4 of [19], $L$ is an isomorphism from the tangent space of $V$ at $w$ to that of $W$ at $(g,\varphi,h)$. Then by the inverse function theorem [5, Section 15], $\Phi$ is an $C^{1}$ diffeomorphism from $B_{\tilde{\rho}}(w)$, a small neighborhood of $w$ in $\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$ to an open neighborhood $U$ of $(g,\varphi,h)$ in $W$. Take $\tilde{\rho}$ small enough, then the functions in $B_{\tilde{\rho}}(w)$ is still uniformly convex in $x$.
@@
Since $g_{\epsilon}=0$ for $0\leq t\leq 2\epsilon$, then $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$ is a solution to equation (3.1) in $\Omega\times(0,2\epsilon]$ and is uniformly convex in $x$. ∎
@@
$u^{(0)}(x)=\varphi(x),\quad u^{(k+1)}(x)=\frac{\partial^{k}}{\partial t^{k}}\Big{|}_{t=0}\Big{(}\det D^{2}u(x,t)-\psi(x,t)\Big{)}.$
Here the right term of the second equality can be recursively expressed by derivatives of $\varphi$ and $\psi$.
The uniqueness of solutions to the equation (2.1) then follows immediately from Lemma 2.1.
@@
Let $\Omega$ be a uniformly convex bounded domain with smooth boundary and $\phi$ be a continuous function in $\overline{Q}_{T}$. Then equation (2.1) admits at most one solution in $\tilde{C}^{2}$ that is convex in $x$.
3.1. Proof of main theorems
@@
Additionally, the higher order regularity of $u$ can also be obtained by classic regularity theory of linear parabolic equations iteratively if $\partial\Omega$ and $\phi$ are smooth enough, and equation (2.1) admits such solution uniquely if the equation satisfies higher order compatibility conditions in Remark 3.2. Specifically, let $k\geq 2$, $0<\beta\leq 1$. Assume that $\partial\Omega$ is of class $C^{k+2,\beta}$, $\phi\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$, and equation (2.1) satisfies the compatibility condition of order $[\frac{k+\beta}{2}]+1$. Then the equation admits a unique solution $u\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$ that is convex in $x$. ∎
3.2. Necessity of condition $(\mathcal{P}3)$
Conditions $({\mathcal{P}1})$, $({\mathcal{P}2})$ and $({\mathcal{P}3})$ were introduced by Ivochkina and Ladyzhenskaia [14] to study (1.1). Note that conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ are natural requirements to ensure $C^{-1}\leq\det D^{2}u\leq C$ and thus guarantee the convexity of $u$. We will take counterexamples to show that condition $({\mathcal{P}3})$ is also necessary for the existence of convex solutions to (2.1).
In one space dimension, take $\Omega=(0,1)$ and $T>1$. Then equation (2.1) reduces to the following initial-boundary value problem
{$&-v_{t}+v_{xx}=\psi(x,t)&&\text{ in }Q_{T}=(0,1)\times(0,T],\\ &v=\phi&&\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $\psi$ and $\phi$ are smooth functions satisfying conditions $({\mathcal{P}1})-({\mathcal{P}3})$. Assume $v$ is a smooth solution to this equation that is convex in $x$. Define a smooth bump function $w:\overline{Q}_{T}\to\mathbb{R}$ by
(3.4) $w(x,t)=\left\{\begin{aligned} &At\cdot e^{\frac{-B}{x^{2}(1-x)^{2}}}&&\text{ in }Q_{T},\\ &0&&\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $A,B$ are positive constants. The function $w$ is non-negative, smooth up to the boundary, and satisfies that
(3.5) $w_{t}=0\ \text{on}\ \partial\Omega\times[0,T];\quad\frac{\partial^{k}w}{\partial x^{k}}=0\ \text{on}\ \partial_{p}Q_{T}\ \ \text{for}\ k=0,1,2,\cdots.$
Set $u=v+w$; then we have
(3.6) {$&-u_{t}+u_{xx}=\psi+(-w_{t}+w_{xx})=:\psi+\rho&&\text{ in }Q_{T},\\ &u=\phi&&\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
Since (3.5) gives that $\rho=0$ on $\partial\Omega\times[0,T]$, then conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ are satisfied for $\phi$ and $\psi+\rho$. Direct computation yields
$w_{t}=Ae^{\frac{-B}{x^{2}(1-x)^{2}}},\quad w_{xx}=-2ABt\cdot e^{\frac{-B}{x^{2}(1-x)^{2}}}\frac{P_{6}(x)}{x^{6}(x-1)^{6}}.$
Here $P_{6}(x)$ is a sixth-degree polynomial given by
$P_{6}(x)=2B-8Bx+(8B-3)x^{2}+16x^{3}-33x^{4}+30x^{5}-10x^{6}.$
One can check that there exist constants $x_{0}\in(0,1/2)$ and $c_{1},c_{2}>0$ depending only on $B$ such that
$\inf_{x\in(0,1)}\rho(x,1)=\rho(x_{0},1)=-A\cdot c_{1}<0,\quad\sup_{x\in(0,1)}\rho(x,1)=\rho(1/2,1)=A\cdot c_{2}>0.$
Then by choosing $A$ sufficiently large, $\rho$ attains both a negative minimum and a positive maximum, each with large absolute values. This destroys the concavity of $\psi+\rho$. For any fixed $t$, the maximum of $w(\cdot,t)$ on $(0,1)$ is $At\cdot e^{-16B}$. Then, the maximum of $w$ can be made arbitrarily large and $u$ fails to remain convex in $x$; therefore, by the comparison principle, equation (3.6) does not admit any solution that is convex in $x$ for all time.
In $n$ space dimension, a counterexample can also be constructed similarly for rotationally symmetric solutions. Let $v(r,t)$ be a smooth radial solution on $(0,1)\times(0,T]$ satisfying
{$&-v_{t}+(v_{r}/r)^{n-1}v_{rr}=\psi(r,t)&&\text{ in }Q_{T}=(0,1)\times(0,T],\\ &v=\phi&&\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
And $\bar{v}(x,t)=v(|x|,t)$ solves the parabolic Monge–Ampère equation
{$&-\bar{v}_{t}+\det D^{2}\bar{v}=\psi(|x|,t)&&\text{ in }U_{T}=B_{1}(0)\times(0,T],\\ &\bar{v}=\phi(|x|,t)&&\text{ on }\partial_{p}U_{T}.\end{aligned}\right.$
We use the same one‐dimensional bump $w(r,t)$ as in (3.4) above and set $\bar{w}(x,t)=w(|x|,t)$. The equation for $\bar{u}=\bar{v}+\bar{w}$ becomes
(3.7) {$&-\bar{u}_{t}+\det D^{2}\bar{u}=\bar{\Psi}(x,t)&&\text{in }U_{T},\\ &\bar{u}=\phi(|x|,t)&&\text{ on }\partial_{p}U_{T},\end{aligned}\right.$
where
$\bar{\Psi}(x,t)=\Psi(|x|,t)=\Psi(r,t)=-v_{t}-w_{t}+\big{(}\frac{v_{r}+w_{r}}{r}\big{)}^{n-1}v_{rr}+\big{(}\frac{v_{r}+w_{r}}{r}\big{)}^{n-1}w_{rr}.$
By (3.5), $\bar{\Psi}$ coincides with $\psi$ on $\partial B_{1}(0)\times[0,T]$, and $\bar{u}$ matches $\bar{v}$ on $\partial_{p}U_{T}$. Thus, conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ hold for arbitrary $A$ and $B$. For sufficiently large $A$, there exist $r_{0}\in(1/2,1)$ and $c_{3},c_{4}>0$ depending on $B$ such that $w_{r}(r_{0},1)=A\cdot c_{3}>0$ and $w_{rr}(r_{0},1)=-A\cdot c_{4}<0$. Since $||v||_{\tilde{C}^{2}(\overline{Q}_{T})}$ is bounded, $\Psi(r_{0},1)$ can be made arbitrarily negative, which destroys the concavity of $\Psi$ in $r$. Additionally, $u$ loses convexity in $r$ when $w$ attains a sufficiently large maximum. From the comparison principle Lemma 2.1, we know that there is no convex solution to equation (3.7).
@@ 4. Gauss curvature flow
(4.1) {$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{\frac{n+1}{2}}}\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\,\;\,\text{on }\partial_{p}Q_{T}.\end{aligned}\right.$
@@
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be convex in $x$ and solve equation (4.1). Suppose that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$, $\partial\Omega\in C^{4}$, $\phi\in\tilde{C}^{4}(\overline{Q}_{T})$, $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$ and $\phi(\cdot,0)$ is uniformly convex. Then we have
@@
(4.2) $\mathcal{L}^{\prime}$ $=-\partial_{t}+\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta}}u^{ij}\partial_{ij}-2\beta\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta+1}}u_{i}\partial_{i}$
@@
(4.3) $\overline{w}=\tilde{\phi}-\frac{1}{q}\log(1+kd_{\partial\Omega})$
@@
Finally, we take $k$ sufficiently large to satisfy
$\log(k\delta)\geq q(\max_{\overline{\Omega}}|\tilde{\phi}|+\max_{\overline{Q}_{T}}|u|),$
then $\overline{w}\leq u$ on $\partial\Omega_{\delta}\times\{t_{0}\}$. Therefore $u\geq\overline{w}$ in $\Omega_{\delta}\times\{t_{0}\}$ by the comparison principle. Then for the exterior normal vector $\nu$ at any point $x_{0}$ on $\partial\Omega$, we have $u_{\nu}\leq w_{\nu}\leq C$. By convexity of $u$ in $x$ and arbitrariness of $t_{0}$, this proves the bound for $|Du|$ on $\overline{Q}_{T}$.
@@
(4.4) $C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u\leq CI_{n-1}\quad\text{on}\quad\partial_{p}Q_{T}.$
@@
(4.5) $|\mathcal{L}^{\prime}\mathcal{T}(u-\phi)|\leq\underline{a}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}$
@@
(4.6) $|\mathcal{T}(u-\phi)|\leq\bar{a}|x^{\prime}|^{2},\quad\text{on $(B_{\varepsilon}\cap\partial\Omega)\times[0,T]$}.$
@@
(4.7) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in $(B_{\varepsilon}\cap\Omega)\times\{0\}$ for all positive constant $A$.}$
@@
(4.8) $\begin{split}\mathcal{L}^{\prime}\zeta&\geq c_{1}\Big{(}\sum_{i=1}^{n}u^{ii}+Mu^{nn}\Big{)}-c_{2}\\ &\geq\frac{1}{2}c_{1}\sum_{i=1}^{n}u^{ii}-c_{2}+c_{3}M^{1/n}.\end{split}$
@@
(4.9) $\mathcal{L}^{\prime}(A\zeta\pm\mathcal{T}(u-\phi))\geq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times(0,T].$
@@
(4.10) $\zeta\leq-\frac{1}{2}\mu\sum_{\beta<n}x_{\beta}^{2}+O(|x^{\prime}|^{3})\leq-\frac{1}{4}\mu|x^{\prime}|^{2},\quad\text{on $(\overline{B_{\varepsilon}}\cap\partial\Omega)\times[0,T]$.}$
@@
(4.11) $\begin{split}\zeta&\leq\frac{1}{2}(B_{\alpha\beta}-\mu\delta_{\alpha\beta})x_{\alpha}x_{\beta}+\frac{1}{2}Mx_{n}^{2}-x_{n}\\ &\leq-\frac{1}{2}\mu|x^{\prime}|^{2}+C|x^{\prime}|^{3}+Mx_{n}^{2}\\ &\leq-c_{4}\varepsilon^{2},\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$
@@
(4.12) $\begin{split}\zeta&\leq C|x^{\prime}|^{2}+Mx_{n}^{2}-x_{n}\\ &\leq(M+C)\varepsilon^{2}-c_{5}\varepsilon\\ &\leq-\frac{1}{2}c_{5}\varepsilon,\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$
@@
(4.13) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{on}\quad\partial(B_{\varepsilon}\cap\Omega)\times[0,T],$
@@
(4.14) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times[0,T].$
@@
(4.16) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{on }\partial_{p}Q_{T}$
@@
Assume $\Omega$ is a smooth and uniformly convex domain in $\mathbb{R}^{n}$, $\phi$ is a smooth function in $\overline{Q}_{T}$ with $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$, and $\phi(\cdot,0)$ is uniformly convex. If equation (4.1) satisfies the compatibility condition of any order, then there exists a unique smooth solution $u$ to the initial boundary value problem (4.1) that is convex in $x$.
@@
Acknowledgement
We would like to thank Professor X.-J. Wang for his suggestions on this topic.